OpenStudy (anonymous):

a radio telescope has a parabolic surface. If the telescope is 1m deep and 12m wide, how far is the focus from the vertex

OpenStudy (anonymous):



OpenStudy (anonymous):

The points (-6,1), (0,0), and (6,1) are on the parabola with the equation of y = ¹ ̷ ₃₆x². Since a = 1/36, then 1/36 = 1/4p, so p = 9 The focus is 9 meters from the vertex. <==ANSWER i hope that helps!! :-)

OpenStudy (anonymous):

Thank you very much!

OpenStudy (anonymous):

no problem